OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false. 4 ⋅ 6 + 5 ⋅ 7 + 6 ⋅ 8 + ... + 4n( 4n + 2) = [4(4n+1)(8n+7) / 6

OpenStudy (anonymous):

@uri @skullpatrol @mathmate

OpenStudy (mathmate):

Have you done mathematical induction?

OpenStudy (anonymous):

I have seen it in my lesson but its extremely vague. Is it simple enough for you to give me a short explanation. Or should I try and figure it out first?

OpenStudy (mathmate):

The three steps are: 1. Verify that the statement is true for n=1 by substitution. 2. \(Assuming\) that the statement is true for case n, show that it is also true for case n+1. This is done by adding the (n+1) term on the left hand side and try to manipulate the expression algebraically so that it becomes the expression on the right, with n replaced by (n+1). 3. Conclude that the statement is true for all n >=1.

OpenStudy (anonymous):

So for step 1 I just plug in 1 for n and solve?

OpenStudy (anonymous):

Or I just solve for n

OpenStudy (mathmate):

You don't really "solve". You put in n=1 to evaluate both sides. Verify that the left hand side equals the right hand side. If they are not equal, check your question, or check your calculations.

OpenStudy (anonymous):

Oh okay! Just a sec while I solve

OpenStudy (mathmate):

Example: show that \(1+2+3+...+n=\frac{n(n+1)}{2}\) Step 1: put n=1, LHS=1, RHS=(1*2)/2=1, So the statement is valid for n=1.

OpenStudy (mathmate):

Step 2: Assume that the statement is true for n, then \(1+2+3+...+n=\frac{n(n+1)}{2}\) for case n+1, we have \(1+2+3+...+n + (n+1)=\frac{n(n+1)}{2}+(n+1)=\frac{n(n+1)+2n+2}{2}=\frac{(n+1)(n+2)}{2}\) Therefore the statement is also true for n+1, hence by mathematical induction, \(1+2+3+...+n=\frac{n(n+1)}{2}\) is true for all n \(\ge\)1.

OpenStudy (anonymous):

Crap... I think you lost me already

OpenStudy (mathmate):

Have you done the first step?

OpenStudy (anonymous):

No.. what are LHS and RHS ?

OpenStudy (mathmate):

It's a mathematical jargon to mean left-hand side and right hand side.

OpenStudy (anonymous):

Okay. And how did you get that equation you had for RHS?

OpenStudy (mathmate):

I think you'd be better off studying your teacher's notes, or study my example to understand how the logic goes before attempting the more difficult problem your teacher gave you.

OpenStudy (mathmate):

For step 1? You'd put n=1 into the LHS and RHS and show that they are equal. If there's no arithmetical error and they are not equal, you just proved that the statement is not true.

OpenStudy (anonymous):

Okay. So unless I messed up I got the left term as 24 and the right as 50

OpenStudy (mathmate):

Let me check.

OpenStudy (mathmate):

That's what I got too, so that means that the statement (that you've written) is not true. Check that there's no typo.

OpenStudy (anonymous):

Looks good to me

OpenStudy (mathmate):

If you look just on the left hand side, 4.6+5.7+.... will not make 4n(4n+2) because then first term is 4.6, second term (n=2) is 8.10, not 5.7 There is definitely a mix-up somewhere.

OpenStudy (anonymous):

Right. So at this point I am done with this problem?

OpenStudy (mathmate):

The problem should read, if the LHS is good: \(4.6+5.7+6.8+...+(n+3)(n+5)=\large \frac{n(2n^2+27n+115)}{6}\)

OpenStudy (mathmate):

I would contact the teacher if there is no typo.

OpenStudy (anonymous):

No the point are multiplication signs

OpenStudy (mathmate):

In any case, study the example I gave you. That would help you when you get the correct problem.

OpenStudy (mathmate):

Yes, I used the points as multiplication signs, just like the original problem.

OpenStudy (anonymous):

And I believe Ive figured out how to complete these problems

OpenStudy (mathmate):

Good! That's the whole point of exercise.

OpenStudy (mathmate):

Good luck with your studies and assignments.

OpenStudy (anonymous):

Awesome! thanks so much!!

OpenStudy (mathmate):

You're welcome!